OpenStudy (anonymous):

Estimate each by rounding to the indicated place value. 14.8557 + 11.7391; Hundereths.

OpenStudy (anonymous):

rounding Hundereths?

OpenStudy (anonymous):

Which number is the hundredths in the first number which number is the hundredths in the second number

OpenStudy (anonymous):

14.8557 + 11.7391=26.5948 rounding Hundereths: 26.5948 =26.60

OpenStudy (anonymous):

I think they want you to round each one and then add them

OpenStudy (anonymous):

14.86 + 11.74 = 26.60

OpenStudy (anonymous):

thank you, Blexting and Nacy Lam.

OpenStudy (anonymous):

oops sorry Nancy I spelled your name wrong.

OpenStudy (anonymous):

if round frist you do like this 14.86 + 11.73=26.59 if continue round : 26.59 = 26.6 No problem

OpenStudy (anonymous):

14.86 + 11.74=26.59 26.6 Sorry I missing